INFRARED LASER VELOCITY MODULATION SPECTROSCOPY OF SiCl^{+} (X^{1}Σ^{+}) UP TO V=6

W. Y. FAN, \underline{Z. LIU} and P. B. DAVIES, Department of Chemistry, Uiversity of Cambridge, Cambridge CB2 1EW, UK.

Many fundamental and hot band absorption lines of ^{28}Si^{35}Cl^{+}, ^{28}Si^{37}Cl^{+}, ^{29}Si^{35}Cl^{+} and ^{30}Si^{35}Cl^{+} have been detected between 630 and 700 cm^{−1} in a SiCl_{4}/He ac discharge using diode laser velocity modulation spectroscopy. The data have been fitted to give mass independent coefficients U_{kl}. The derived spectroscopic constants for ^{28}Si^{35}Cl^{+} include ω_{e} = 678.24316(18) cm^{−1} and B_{e} = 0.2870288(14) cm^{−1}. The equilibrium internuclear distance is r_{e}=1.9439105(46)Å. The equilibrium dissociation energy calculated from the expanded Dunham-Morse potential is D_{e}=49431(57)cm^{−1}.